Assume you have computed an analysis of variance for three groups with 10 observations per group. The degrees of freedom for the F test are:

Group of answer choices

2 and 27

3 and 30

2 and 30.

2 and 10.
Answer #1

Groups = k = 3

n = 3 * 10 = 30

Degrees of freedom = (k-1) , (n-k)

= ( 3 - 1) , (30 - 3)

= 2 and 27